# Conformalized Strategy-Proof Auctions

### v4

@@ Abstract
Auctions are key for maximizing sellers’ revenue and ensuring truthful bidding among buyers. Recently, an approach known as differentiable economics based on machine learning (ML) has shown promise in learning powerful auction mechanisms for multiple items and participants. However, this approach has no guarantee of strategy-proofness at test time. Strategy-proofness is crucial as it ensures that buyers are incentivized to bid their true valuations, leading to optimal and fair auction outcomes without the risk of manipulation. In this work, we propose a formulation of statistical strategy-proofness auction mechanism, ensuring that the probability of regret exceeding a predefined threshold is strictly controlled. Building upon conformal prediction techniques, we develop an auction acceptance rule that leverages regret predictions to guarantee that the data-driven auction mechanism meets the statistical strategy-proofness requirement with high probability. Our approach represents a practical middle-ground between two extremes: forcing zero-regret at the cost of significant revenue loss, and naively using ML to construct auctions with the hope of attaining low regret at test time. Numerical experiments demonstrate the necessity of the proposed method, the validity of our theoretical result, and its applicability.
@@ 1 Introduction
Auction design represents a fundamental component of economic theory, holding significant practical relevance in various industries and the public sector for organizing the sale of products and services [Milgrom, 2004, Krishna, 2009]. Auctions serve as a crucial mechanism for allocating scarce goods and generate billions in revenue across domains such as online advertising, sourcing, and more [Cramton, 1997, Sandholm, 2007, Leyton-Brown et al., 2017, Milgrom, 2017, Roth, 2018]. Notable examples include sponsored search auctions by search engines such as Google, and auctions on platforms such as eBay [Edelman et al., 2007]. Within the standard independent private valuations model, each bidder possesses a valuation function—the true values they are willing to pay—over subsets of items, independently drawn from possibly distinct distributions. The auctioneer is presumed to have knowledge of these value distributions, a crucial factor in auction design. However, a primary challenge arises from the private nature of valuations, as bidders might not disclose their true valuations, complicating the auction process.
1.1 Data-Driven Auctions
The pursuit of optimal auction mechanisms has been a subject of great interest to economics for many years. Myerson auction’s [Myerson, 1981] stands out as the optimal mechanism for auctioning a single item to several bidders. In contexts where multiple items are auctioned to a single bidder, the concept of optimality has been largely delineated [Manelli and Vincent, 2006, Pavlov, 2011, Daskalakis et al., 2015, Kash and Frongillo, 2016]. However, when it comes to auctioning multiple items to various bidders, optimal mechanisms have been identified only in simplified scenarios, such as those discussed by Yao [2017]. Indeed, the existence of an optimal mechanism in the multi-bidder multi-items setting remains an open question even in simple scenario such as two or more bidders and several items.
The predictive power of deep neural network (DNN) models, coupled with the quest for analytical outcomes, drive the research by Dütting et al. [2019] to a data-driven solution; see also [Dütting et al., 2024]. While this data-driven approach is not guaranteed to maximize revenue or be entirely strategy-proof, empirical evidence shows that the resulting deep-learning solutions are close to recapturing the outcome of well-known optimal auction designs, when they exist.
1.2 The Lack of Strategy-proofness in Data-Driven Auctions
Strategy-proofness relates closely to the notion of regret in auctions, a crucial metric that indicates whether the outcomes of a given auction mechanism are strategy-proof. Specifically, a mechanism with low regret ensures that bidders have little incentive to deviate from their true valuations. In other words, a low-regret mechanism encourages truthful bidding, which is a cornerstone of strategy-proofness [Dütting et al., 2015, Milgrom, 2004, Myerson, 1981]. As such, the importance of low regret in auctions cannot be overstated. In a practical sense, low regret translates to higher revenue for the auctioneer and more efficient allocation of resources among bidders. Moreover, it fosters trust in the auction mechanism, as participants can be confident that truthful bidding is in their best interest [Krishna, 2009].
@@ 1.2 The Lack of Strategy-proofness in Data-Driven Auctions
A critical issue in deploying data-driven mechanisms for auctions, particularly those based on DNNs, is the lack of statistical guarantees during test time for new auctions. While these models are trained to minimize regret on average, leading to low-regret outcomes in many cases, they occasionally produce predictions for allocations and payments that grossly violate the low-regret requirement at test time. This limitation is visualized in Figure 1, illustrating that a well-known data-driven auction mechanism—RegretNet by Dütting et al. [2019]—fails to attain low-regret levels. While the average regret across many scenarios tends to be low, the presence of high regret values in the tail of the distribution is evident in the red zone in Figure 1. We stress this point in Figure 2, which shows an increase in the maximum regret obtained at test time as the number of auctions evaluated grows, suggesting that the model’s reliability diminishes as it encounters a wider variety of auction contexts.
@@
Figure 1: Empirical distribution of actual regret values of RegretNet model, evaluated on 10,000 test auctions; the auction setup is $2\times 3$. For additional insights into the experimental configuration of this model, please refer to Section 4.
@@
Figure 2: Maximum regret value obtained by RegretNet model as a function of the number of test auctions. The auction setup is $2\times 3$.
@@
The discussion above reveals a significant limitation of data-driven auction design: there is no assurance that the regret will remain low with any level of confidence once the model is deployed in real-world scenarios. This undesired behaviour could result in substantial revenue losses for the auctioneer and unfair outcomes for the bidders, undermining the trust and efficiency of the auction system. The primary goal of this paper is to tackle this challenge by selectively accepting auctions for which the unknown test-time regret is unlikely to exceed a desired maximum regret level. However, before presenting our proposal, we first pause to review related work.
@@ 1.3 Related Work
Given the theoretical and practical importance of strategy-proofness, there has been growing activity in enhancing the reliability of data-driven auctions. Curry et al. [2020] modify the RegretNet model [Dütting et al., 2019] to enable certifiable strategy-proofness using mixed integer programming. However, the certification procedure imposes strict constraints both on the architecture design and the number of trainable parameters—as it involves solving a large mixed integer programming problem—limiting the model’s performance. By contrast, as we will see, the framework we propose in this paper can be applied to virtually any auction model, regardless of its complexity and architectural design.
A different line of work by Peri et al. [2021], Rahme et al. [2021b, a], Ivanov et al. [2022] offers novel learning approaches and architectures to formulate actions with low average regret. This stands in striking contrast with our method, which achieves a high probability bound on the worst-case, maximal regret obtained at test time.
Another line of work for forming reliable data-driven auctions builds on the parameterization of allocation menus that are guaranteed to be strategy-proof. This area was recently advanced by Curry et al. [2023] and Duan et al. [2023], who optimize the parameters of affine maximizer auctions (AMA) [Roberts, 1979]. While AMA-based approaches are strategy-proof, they often yield lower revenue than standard DNN-based methods. Indeed, in Section 4, we show that the approach by [Duan et al., 2023] results in a revenue loss of about 45% compared to our method. This emphasizes the hardness of attaining exact strategy-proof auctions while constructing data-driven auctions with high revenue.
1.4 Overview and Contributions
The challenge of forming zero-regret, strategy-proof auctions with high revenue drives us to propose a new, statistical approach for constructing reliable data-driven auctions. Our key contribution is the introduction of an auction acceptance rule that selectively rejects predicted auctions whose regret exceeds a pre-defined threshold. In turn, the advantage of our approach is that it guarantees, with high probability (e.g., 99%), that accepted auctions at test time do not exceed a pre-defined max regret value (e.g., of level 0.05).
@@ 1.4 Overview and Contributions
As such, our work introduces a new practical paradigm for balancing the trade-off between strategy-proofness and revenue discussed above. Our approach can be seen as a synthesis of two extremes: (i) enforcing zero regret at the expense of significant revenue loss, and (ii) naively using machine learning to design auctions with the hope of achieving low regret at test time. Indeed, our experiments show that the accepted auctions obtained by our approach are of low regret while having revenue nearly the same as the baseline auction model by Dütting et al. [2019].
@@ 2.1 Problem Setup
We consider an auction environment, comprising of a set of $n$ bidders, denoted by $[n]=\{1,\ldots,n\}$, and a set of $m$ items, denoted by $[m]=\{1,\ldots,m\}$. Each bidder $i\in[n]$ possesses a private valuation $v^{i}\in\mathbb{R}^{m}$ which maps item sets $S\subseteq[m]$ to their value $v^{i}(S)\in\mathbb{R}_{\geq 0}$ for the bidder. The valuation matrix $v=[v^{1},\ldots,v^{n}]$ is assumed to be drawn from a distribution $P$.
In an auction, participants submit bids that may not necessarily represent their true valuations. The auction mechanism, denoted as $f(b)=[g(b),p(b)],$ comprises an allocation rule $g(b)$ and a payment rule $p(b)$. Specifically, the allocation rule is defined as $g:\mathbb{R}_{\geq 0}^{n\times m}\to[0,1]^{n\times m},$ determining how the $m$ items are assigned to the $n$ bidders. The payment rule is given by $p:\mathbb{R}_{\geq 0}^{n\times m}\to\mathbb{R}_{\geq 0}^{n\times m},$ specifying the amount each bidder must pay. Here, the bid profile of all bidders is represented as $b\in\mathbb{R}_{\geq 0}^{n\times m}$. The primary goal of the auction mechanism is to allocate the items efficiently while setting payments, often with the objective of maximizing revenue.
To effectively frame strategy-proof auctions as a learning task, we shall design a DNN trained to minimize the regret while maximizing the bidders’ utility. To this end, we denote the trained auction mechanism by ${f}_{\hat{\theta}}(b)=[{g}_{\hat{\theta}}(b),{p}_{\hat{\theta}}(b)]$, where the vector $\theta$ represents trainable parameters. Here, the ‘hat’ symbol over the parameters $\hat{\theta}$ indicates that these are fixed, i.e., the model is pre-trained. Using the notations from [Dütting et al., 2019], the utility of bidder $i$, denoted by $u^{i}_{\hat{\theta}}(v^{i};b)$, is defined as the difference between their valuation of the allocated items and the payment made, i.e.,
@@
Finally, the formulation of the learning task of an auction setup with valuations drawn from $P$ is as follows:
$\min_{\theta\in\mathbb{R}^{d}}$ $\quad\underset{v\sim P}{\mathbb{E}}\left[-\sum_{i\in N}p_{\theta}^{i}(v)\right]$ [REVENUE] (3)
s.t. $\operatorname{\textbf{rgt}}_{\theta}^{i}(v;b)=0,\hskip 7.39755pt\ \forall i\in[n]$  [REGRET]
$u_{\theta}^{i}(v^{i};b)\geq 0,\hskip 15.6491pt\forall i\in[n]$  [IR]
[…]
In plain words, the goal of this optimization problem is to maximize revenue while minimizing regret, promoting incentive rationality (IR) and a feasible mechanism.
@@ 2.2 Conformal Prediction
Conformal prediction [Papadopoulos et al., 2002, Vovk et al., 2005, Lei and Wasserman, 2014] offers a generic approach, applicable to any predictive model, for quantifying prediction uncertainty. In recent years, conformal prediction has increasingly gained traction across various domains, demonstrating its broad utility and effectiveness in practical settings [Angelopoulos and Bates, 2023]. This method utilizes a subset of holdout calibration data to quantify the prediction error of the model. These errors are then used to construct a range of plausible predictions—a prediction interval—that covers the unknown test outcome with high probability.
In what follows, we describe in more detail how to construct prediction intervals for regression problems, as this task is tightly connected to our goal of reliably predicting unknown regret values at test time. Let $x\in\mathcal{X}$ represent the input features to the model, $y\in\mathcal{Y}$ denote the true output, and ${f}_{\hat{\theta}}(x)$ be the predicted output generated by the model ${f}_{\hat{\theta}}$ parametrized by ${\hat{\theta}}$. The process of conformal prediction for general inputs and outputs can be outlined as follows. First, we apply ${f}_{\hat{\theta}}$ on the labeled calibration data $\mathcal{D}_{\textup{cal}}=\{(x_{k},y_{k})\}_{k=1}^{K}$ to construct a heuristic notion of prediction error for each calibration point $(x_{k},y_{k})$. We refer to this notion of error as a nonconformity score, denoted by $s(x,y)\in\mathbb{R}$. The next step involves calculating $\hat{q}_{\alpha}$, the $(1-\alpha)(1+1/K)$ empirical quantile of the calibration scores $s(x_{k},y_{k})$, $k=1,\dots,K$. This quantile is used to construct a prediction interval for a new test point $x_{\textup{test}}$, which includes all the output values $\tilde{y}\in\mathcal{Y}$ whose corresponding nonconformity scores fall below the quantile $\hat{q}_{\alpha}$, i.e.
@@
The key property of $C(x_{\textup{test}})$ is that it is guaranteed to cover the unknown test label $y_{\textup{test}}$ at the desired level of coverage $1-\alpha$, as formally stated in Theorem 1.
Theorem 1 (Conformal Coverage Guarantee; Vovk, Gammerman, and Saunders 1999).
@@
Notably, the standard conformal prediction approach cannot be applied “as is” in auction settings. This is because the regret value, which we aim to control, depends on the user’s bids $b$, valuations $v$, and the allocation and payment determined by the auction model ${f}_{\hat{\theta}}$. In other words, the “label” $y$ is now a complex function of both the model and the data, requiring a customized adaptation of conformal prediction to fit this auction context—this is the focus of our work.

### v5

# Statistically Truthful Auctions via Acceptance Rule

@@ Abstract
Auctions are key for maximizing sellers’ revenue and ensuring truthful bidding among buyers. Recently, an approach known as differentiable economics based on machine learning (ML) has shown promise in learning powerful auction mechanisms for multiple items and participants. However, this approach has no guarantee of strategy-proofness at test time. Strategy-proofness is crucial as it ensures that buyers are incentivized to bid their true valuations, leading to optimal and fair auction outcomes without the risk of manipulation. In this work, we propose a formulation of statistical strategy-proofness for auction mechanisms. Specifically, we offer a method that bounds the regret—quantifying deviation from truthful bidding—below a pre-specified level with high probability. Building upon conformal prediction techniques, we develop an auction acceptance rule that leverages regret predictions to guarantee that the data-driven auction mechanism meets the statistical strategy-proofness requirement with high probability. Our method—Statistically Truthful Auctions via Acceptance Rule (SPACE )—represents a practical middle-ground between two extremes: enforcing truthfulness—zero-regret—at the cost of significant revenue loss, and naively using ML to construct auctions with the hope of attaining low regret, with no test-time guarantees.
Figure 1: Overview of Statistically Proofed Auctions via Conformal Estimation (SPACE ). Bidders submit bids for multiple items, forming a bid matrix. An auction model (mechanism) processes these bids to determine allocation and payment. A regret model estimates the incentive to deviate from truthful bidding. Using regret predictions and holdout calibration data, a selective acceptance rule ensures statistical strategy-proofness, accepting only outcomes with low estimated regret. This balances revenue and incentive compatibility.
@@ 1 Introduction
Auctions are a core tool of economic theory and practice, allocating valuable goods and generating billions across markets such as online advertising and procurement; notable implementations include sponsored search (e.g., Google) and eBay auctions [22, 16, 5, 32, 18, 20, 31, 13]. In the standard independent private valuations model, each bidder has a valuation function over subsets of items drawn independently, possibly from distinct distributions, the auctioneer knows these distributions, yet valuations remain private—creating incentives to misreport and complicating design.
Optimal mechanisms are known only in several special cases such as Myerson’s setting—single-item multi-bidder auctions [23]. For multiple items with a single bidder, much of the landscape is characterized [19, 25, 8, 15]. For multiple items and multiple bidders, optimal mechanisms are known only in simplified settings [40], and their characterization remains open even with a few bidders and several items. Leveraging the predictive power of deep neural networks (DNN), Dütting et al. [11] and follow-ups [12, 9, 27, 28, 14, 26, 7, 39, 33] propose data-driven mechanisms that, while lacking guarantees of revenue optimality or full strategy-proofness, empirically approximate known optimal designs when these are known.
1.1 The Challenge of Strategy-Proofness in Data-Driven Auctions
Strategy-proofness is tightly linked to regret in auctions: regret measures the incentive of bidders to deviate from the truth, so a mechanism with low regret leaves bidders little reason to misreport their true valuations. Put differently, low regret promotes truthful bidding, yielding at the limit a strategy-proof auction mechanism [10, 22, 23]. The practical stakes are clear: lower regret supports greater trust in the mechanism, since participants can be confident that bidding truthfully is roughly in their best interest, and thus enables the auctioneer to guarantee higher revenue and more efficient allocation of resources across bidders [16].
@@ 1.1 The Challenge of Strategy-Proofness in Data-Driven Auctions
A critical deployment issue for many data-driven—DNN-based—mechanisms is the absence of statistical guarantees at test time for new auctions. Although these models are trained to minimize regret on average and often achieve low-regret outcomes, they can still output allocations and payments that severely violate the low-regret requirement on unseen instances. This limitation is illustrated in Figure 2(a): the well-known RegretNet mechanism [11] does not consistently attain low regret, with a visible tail of high-regret outcomes (the red zone). While average regret may be small across many scenarios, the tail behavior persists. Figure 2(b) emphasizes this point by showing that the maximum test-time regret rises as more auctions are evaluated, suggesting diminished reliability as the model encounters a broader range of auction contexts.
@@
(a) Empirical distribution of actual regret values of RegretNet model, evaluated on 10,000 test auctions.
@@
(b) Maximum regret value obtained by RegretNet model as a function of the number of test auctions.
Figure 2: Regret behavior of RegretNet in the 22-bidder, 33-item auction setting. For additional details on the experimental configuration of this model, please refer to Section 4.
@@
The discussion above reveals a significant limitation of cutting-edge data-driven auction design: there is no assurance that the regret will remain low with any level of confidence once the model is deployed in real-world scenarios. This undesired behavior undermines the trustworthiness of the auction system, and could lead to inefficiency, substantial revenue losses for the auctioneer, and unfair outcomes for the bidders. The primary goal of this paper is to tackle this challenge by selectively accepting auctions for which the unknown test-time regret is unlikely to exceed a desired maximum regret level.
@@ 1.2 Related Work
Given the theoretical and practical importance of strategy-proofness, there has been growing activity in enhancing the reliability of data-driven auctions. Curry et al. [6] modify the model proposed by [11] to enable certifiable strategy-proofness using mixed integer programming. However, the certification procedure imposes strict constraints both on the architecture design and the number of trainable parameters—as it involves solving a large mixed integer programming problem—limiting the model’s performance. By contrast, as we will see, the framework we propose in this paper can be applied to virtually any auction model, regardless of its complexity and architectural design.
A different line of work by Peri et al. [26], Rahme et al. [28, 27], Ivanov et al. [14] offers novel learning approaches and architectures to formulate actions with low average regret. This stands in striking contrast with our method, which achieves a high probability bound on the worst-case, maximal regret obtained at test time.
Another line of work for forming reliable data-driven auctions builds on the parameterization of allocation menus that are guaranteed to be strategy-proof. This area was recently advanced by Curry et al. [7] and Duan et al. [9], who optimize the parameters of affine maximizer auctions (AMA) [29]. While AMA-based approaches are strategy-proof, they often yield lower revenue than standard DNN-based methods. Indeed, in Section 4, we show that the approach by [9] results in a revenue loss of about 45% compared to our method. This emphasizes the hardness of attaining exact strategy-proof auctions while constructing data-driven auctions with high revenue.
1.3 Overview of Contributions
The challenge of forming guaranteed low-regret auctions with high revenue drives us to propose a new, statistical approach for constructing reliable data-driven auctions. Our key contribution is the introduction of an auction acceptance rule that selectively rejects predicted auctions whose regret exceeds a pre-defined threshold. In turn, the advantage of our approach is that it guarantees, with high probability (e.g., 99%), that accepted auctions at test time do not exceed a pre-defined max regret value (e.g., of level 0.05).
@@ 1.3 Overview of Contributions
As such, our work introduces a new practical paradigm for balancing the trade-off between strategy-proofness and revenue discussed above. Our approach can be seen as a synthesis of two extremes: (i) enforcing zero regret at the expense of significant revenue loss, and (ii) naively using machine learning to design auctions with the hope of achieving low regret at test time. Indeed, our experiments show that the accepted auctions obtained by our approach are of low regret while having revenue nearly the same as the baseline auction model by [11]. As illustrated in Figure 3, our statistically strategy-proof mechanism achieves a favorable balance between revenue and strategy-proofness, outperforming traditional strategy-proof models in terms of revenue while maintaining controlled regret.
Figure 3: Comparison of revenue across different auction mechanisms. Strategy-proof methods (green) ensure truthful bidding, but yield lower revenue. Non-strategy-proof methods (red) achieve higher revenue at the cost of truthfulness. Our proposed statistically strategy-proof mechanism (blue) balances revenue and strategy-proofness. See Table 2 for references.
@@ 2.1 Problem Setup
We consider an auction environment, comprising of a set of $n$ bidders, denoted by $n]=\{1,\ldots,n\}$, and a set of $m$ items, denoted by $m]=\{1,\ldots,m\}$. Each bidder $i\in[n]$ possesses a private valuation $v^{i}\in\mathbb{R}^{m}$ which maps item sets $S\subseteq[m]$ to their value $v^{i}(S)\in\mathbb{R}_{\geq 0}$ for the bidder. The valuation matrix $v=[v^{1},\ldots,v^{n}]$ is assumed to be drawn from a distribution $P$.
In an auction, participants submit bids $b^{i}\in\mathbb{R}^{m}_{\geq 0}$ that may not necessarily represent their true valuations. The bid profile of all bidders is represented by $b\in\mathbb{R}_{\geq 0}^{n\times m}$. The auction mechanism, denoted by $f(b)=[g(b),p(b)],$ comprises of an allocation rule $g(b)$ and a payment rule $p(b)$. Specifically, the allocation rule is defined as $g:\mathbb{R}_{\geq 0}^{n\times m}\to[0,1]^{n\times m},$ determining how the $m$ items are assigned to the $n$ bidders. The payment rule is given by $p:\mathbb{R}_{\geq 0}^{n\times m}\to\mathbb{R}_{\geq 0}^{n\times m},$ specifying the amount each bidder must pay for each item. The primary goal of the auction mechanism is to allocate the items while setting payments, often with the objective of maximizing revenue.
To effectively frame strategy-proof auctions as a learning task, we shall design a DNN trained to minimize the regret while maximizing the bidders’ utility. To this end, we denote the trained auction mechanism by ${f}_{\hat{\theta}}(b)=[{g}_{\hat{\theta}}(b),{p}_{\hat{\theta}}(b)]$, where the vector $\theta$ represents trainable parameters. Here, the ‘hat’ symbol over the parameters $\hat{\theta}$ indicates that these are fixed, i.e., the model is pre-trained. Using the notations from [11], the utility of bidder $i$, denoted by $u^{i}_{\hat{\theta}}(v^{i};b)$, is defined as the difference between their valuation of the allocated items and the payment made, i.e.,
@@
With valuations function at hand, choose mechanism parameters—train a DNN model—to maximize expected revenue, minimize the regret (strategy-proofness), subject to individual rationality, and feasibility. For further details, see Dütting et al. [11].
A notable feature of data-driven auctions is the prevalence of rejection, where the auction mechanism chooses not to allocate any items—formally denoted as allocation $A_{0}$ [3, 21, 34]. Such no-allocation outcomes are common in practice, particularly in online platforms like ad auctions or marketplaces, where not every auction leads to a match between supply and demand. Rejection serves multiple purposes: it avoids inefficient allocations when the bids do not meet a reserve price or quality threshold, preserves fairness among participants, and can be used strategically to maintain long-term revenue or user experience.
@@ 2.2 Conformal Prediction
Conformal prediction [24, 38, 17] offers a generic approach, applicable to any predictive model, for quantifying prediction uncertainty. In recent years, conformal prediction has increasingly gained traction across various domains, demonstrating its broad utility and effectiveness in practical settings [1]. This method utilizes a subset of holdout calibration data to quantify the prediction error of the model. These errors are then used to construct a range of plausible predictions—a prediction interval—that covers the unknown test outcome with high probability.
In what follows, we describe in more detail how to construct prediction intervals for regression problems, as this task is tightly connected to our goal of reliably predicting unknown regret values at test time. Let $x\in\mathcal{X}$ represent the input features to the model, $y\in\mathcal{Y}$ denote the true output, and ${f}_{\hat{\theta}}(x)$ be the predicted output generated by the model $^{f}_{\hat{\theta}}$ parametrized by $^{\hat{\theta}}$. The process of conformal prediction for general inputs and outputs can be outlined as follows. First, we apply $^{f}_{\hat{\theta}}$ on the labeled calibration data $\mathcal{D}_{\textup{cal}}=\{(x_{k},y_{k})\}_{k=1}^{K}$ to construct a heuristic notion of prediction error for each calibration point $(x_{k},y_{k})$. We refer to this notion of error as a nonconformity score, denoted by $s(x,y)\in\mathbb{R}$. The next step involves calculating $\hat{q}_{\alpha}$, the $(1-\alpha)(1+1/K)$ empirical quantile of the calibration scores $s(x_{k},y_{k})$, $k=1,\dots,K$. This quantile is used to construct a prediction interval for a new test point $x_{\textup{test}}$, which includes all the output values $\tilde{y}\in\mathcal{Y}$ whose corresponding nonconformity scores fall below the quantile $\hat{q}_{\alpha}$, i.e.
@@
The key property of $C(x_{\textup{test}})$ is that it is guaranteed to cover the unknown test label $y_{\textup{test}}$ at the desired level of coverage $1-\alpha$, as formally stated in Theorem 2.1.
Theorem 2.1 (Conformal Coverage Guarantee; Vovk, Gammerman, and Saunders 1999).
@@
Notably, the standard conformal prediction approach cannot be applied “as is” in auction settings. This is because the regret value, which we aim to control, depends on the user’s bids $b$, valuations $v$, and the allocation and payment determined by the auction model $^{f}_{\hat{\theta}}$. In other words, the “label” $y$ is now a complex function of both the model and the data, requiring a customized adaptation of conformal prediction to fit this auction context—this is the focus of our work.